# Integrated Radiation-Magneto-Hydrodynamic Simulations of Magnetized Burning Plasmas. I. Magnetizing Ignition-Class Designs

### v1

@@ Abstract
Motivated by breakthroughs in inertial confinement fusion (ICF), first achieving ignition conditions in National Ignition Facility (NIF) shot N210808 and then laser energy breakeven in N221204,abu1; abu2 modeling efforts here investigate the effect of imposed magnetic fields on integrated hohlraum simulations of igniting systems. Previous NIF experiments have shown fusion yield and hotspot temperature to increase in magnetized, gas-filled capsules moody1 in line with scalings.walsh1 In this work, we use the 2D radiation-magnetohydrodynamics code Lasnexzimmerman75 with a Livermore ICF common model.strozzi1 Simulations are tuned to closely approximate data from unmagnetized experiments. Investigated here is the effect of imposed axial fields of up to 100 T on the fusion output of historically high-performing ICF shots, specifically the record BigFoot shot N180128, and HYBRID-E shots N210808 and N221204. The main observed effect is an increase in the hotspot temperature due to magnetic insulation. Namely, electron heat flow is constrained perpendicular to the magnetic field and alpha trajectories transition to gyro-orbits, enhancing energy deposition into the hotspot and cold fuel.perkins2 In addition, we investigate the impact of applied magnetic fields to future NIF designs, specifically an example Enhanced Yield Capability design with 3 MJ of laser energy as well as a high-, low implosion velocity "Pushered Single Shell" design. In conclusion, magnetization with field strengths of 5-75 T is found to increase the burn-averaged ion temperature by 50% and the neutron yield by 2-12. Specifically, we see notable yield enhancement of at least 50% with only a 5-10 T applied magnetic field for N221204, while a 65 T field on N210808 with drive symmetrization gives an 8 increase in yield. This is all without further design optimization to best take advantage of an applied B field, which promises even greater improvements for future designs tailored specifically towards magnetization.
@@ I Introduction
Magnetized ICF is a relatively old concept, dating back at least to the 1980’s in established literature.lindemuth; jones; guskov The basic promise is that of boosted ignition conditions and yields via reduced conduction losses and enhanced alpha energy deposition into the ignited hotspot due to enhanced magnetic field strengths, where Tesla-scale applied magnetic fields are compressed by the implosion. In Ref. lindemuth, reduced order models informed a relatively large parameter scan suggesting magnetization, starting from an already compressed capsule, could enable designs with less stringent fuel densities, lower implosion velocities, and lower drive requirements that also have significant fuel burn-up, a motivation that continues in recent studies.ho1 Ref. jones suggests that magnetization can enhance the yield for volumetric burn, particularly when burn propagation does not inherently occur but may be detrimental in robustly burning circumstances. This is the origin of the concept that magnetic fields may inhibit burn propagation in igniting targets. One of the primary conclusions of the current study is that applied magnetic fields do not inherently degrade igniting targets and, in fact, may more than double the yield of such designs.
@@
Experiments have subsequently validated several of the basic ideas of these early modeling studies, with the notable exception that no magnetized, cryogenic implosions with ice layers have been conducted thus far. Cylindrical and spherical implosions with direct laser drive at the OMEGA laser facility knauer1; chang; hohenberger demonstrated field compression and magnetized hotspot insulation according to radiation-magnetohydrodynamic (rad-MHD) modeling. Pulsed-power-driven implosions of magnetized cylindrical liners at the Sandia Z Machine provide similar validation in a different regime, relevant to the MagLIF fusion concept.slutz; gomez More recent hohlraum-driven NIF experiments moody1; sio1; strozzi1 with an imposed initial field up to 28 T have shown yield increases up to 2.9x and burn-averaged increases up to 1.4x, consistent with 2D Lasnex rad-MHD modeling.strozzi1 These experiments used gas-filled capsules and 1 MJ of laser energy, due to current practical constraints that magnetized NIF shots must be done at room temperature and with low enough laser energy to avoid risk of optics damage from Brillouin backscatter. Magnetized NIF experiments also examine other topics, like laser-heated gas pipes relevant to MagLIFpollock and high temperature X-ray sources.kemp; leal
@@
This study is the first to present state-of-the-art, fully-integrated rad-MHD modeling of hohlraum-driven, ignition-class implosions with an imposed magnetic field. Where early modeling studies were restricted to reduced-order rad-MHD calculations with strictly azimuthal fields or capsule-only simulations with a frequency-dependent-source (FDS) drive, we are now able to rigorously model magnetized ICF at an appreciable level of fidelity. Modern rad-MHD codes can model full 2D and 3D capsule and hohlraum systems with expanded, non-ideal MHD. These have found that, with the application of more realistic axial fields, initially 10 T fields can be compressed to 10 kT levels to enhance ICF designs, due to factors such as increased robustness to surface roughness, low-mode shape, and residual kinetic energy.perkins1; perkins2; strozzi2; ho2 3D modeling with the Gorgon code of the planar magnetized Rayleigh-Taylor instability has shown reduced growth for modes with k-vectors parallel to the field,walsh1 though magnetized instability modeling of a spherical, integrated implosion has not been published thus far. In addition, there is work looking into the effect of magnetization on hot electron dynamicsstrozzi2 and burn propagation, where some inhibition of burn across field lines has been observed in ignition-style designs.oneill
@@
In this study we model various hohlraum-driven, ignition-class NIF designs and the effect magnetization has on them. It is found that we are able to routinely enhance the yield of full-scale shots like N180128, N210808, and N221204 by at least a factor of while potentially approaching for N210808 at 65 T. The shot nomenclature stands for NIF shot , year , month , and day . There is a notable increase in all three cases with the application of just a 5-10 T magnetic field. The currently studied Enhanced Yield Capability (EYC) design very robustly ignites,maclaren1 meaning the benefit of magnetization is significantly reduced at least for the ad-hoc application of an external field. The Pushered Single Shell (PSS) suggests a way to better take advantage of magnetization given its cooler baseline temperature.maclaren2; maclaren3; dewald; ho3
@@ II NIF Designs for Magnetization
In this study, we focused on five NIF designs for magnetization, three historically high-performing shots, and two potential future designs. The design parameters can be found in Table 1. The selection of shots was not arbitrary; they were all high-performing designs either of historical importance or notable promise based on best practices. All targets used a deuterium-tritium (DT) ice fuel layer and a high-density carbon ablator (HDC) doped with tungsten (W), except for PSS, which used a beryllium (Be) ablator doped with molybdenum (Mo) and traces of argon (Ar) from production. The PSS design is even more unique in that at the outer radius it begins with just Be but then smoothly grades up in density using a dopant. is the difference in wavelength between the inner (23∘, 30∘) and outer (44∘, 50∘) laser cones, which is known as two-color tuning. However, for EYC and PSS we used three- and four-color tuning as well, i.e., separate for 23-30∘ and 44-50∘ quad pairs, to help with shape and reduce potential stimulated Brilloiun scattering (SBS). In this paper we will define in terms of a triplet, where the first number defines two-color wavelength difference (inners versus outers), the second the three-color (23-30∘), and the third the four-color (44-50∘), i.e., but . It is acknowledged that there are many other designs that could be explored, the potential parameter space is enormous. Alternative approaches that could be integrated into a magnetized ICF design were not considered, such as plastic ablators, fast and shock ignition,strozzi3 and wetted foams, for example.
@@
The first shot in consideration was N180128. This record BigFoot shot achieved a yield of kJ and keV baker and was the record yield at the time. Besides its historical importance, one intention for investigating this design was to see if magnetization could take a high-performing, non-igniting design and push it over the so-called ignition cliff. This was not found to be the case for N180128 with no other changes but may be possible with further optimization.walsh4 The second was N210808,abu1 a HYBRID-E (High Yield Big Radius Implosion Design, the E-th iteration) design, which was the first to achieve greater than 1 MJ yield and exceed all extant Lawson criteria for ICF ( MJ, keV). The third case example was N221204,abu2 the first shot to achieve energy breakeven, where more energy was generated than delivered to the hohlraum ( MJ, keV). There have been many minor, cumulative modifications in designs with each shot, but the most pronounced change had been in the increase in laser energy ( MJ, MJ, MJ), and notably the greater radiation temperature efficiency of the Hybrid-E hohlraum.
@@
Two future designs were considered as well. Specifically, a preliminary EYC design,maclaren1 that extends standard HYBRID-E designs from 2 MJ to 3 MJ of laser energy with a peak power of 450 TW. Simple hydroscaling of a design like N221204 is not an option due to laser optics damage exacerbated by higher laser powers. Therefore, new design parameter spaces must be explored. The second is the PSS design driven by 1.9 MJ of laser energy,maclaren2; dewald; ho3 which is a platform tailored to radiochemistry studies and uses a density gradient of Mo-doped Be as the ablator instead of CH or HDC. This design is perhaps ideal for magnetization because it inherently has a slower, high , low-temperature implosion that can be directly enhanced by the application of an external magnetic field.
@@
In Fig. 1 we show a depiction of the hohlraum and target capsules relative to the experimental properties in Table 1. Four of the shots had qualitatively similar HDC capsule designs, while PSS has a unique design. A comparison of the laser power profiles and corresponding hohlraum radiation temperature time history can be seen in Figs. 2 (a) and (b), where is the average radiation temperature across a spherical annulus outside the ablation front of the capsule. N180128 has a distinct BigFoot design while N210808, N221204, and EYC all have HYBRID-E designs. PSS has a distinctly longer foot because Be does not need a picket, which is deleterious to symmetry control, due its much lower melt pressure that keeps the 1st shock rarefaction clear of re-freeze and better ablation front stability.zylstra2018 HDC on the other hand uses a picket due to the higher melt temperature. As is apparent, HYBRID-E designs presented a more efficient approach than BigFoot while EYC has an extended peak to take advantage of greater energy without the higher peak power dangerous to the optics. In Fig. 2 (b) we have raised to the fourth power and normalized to a standard value of (300 eV)4 to represent x-ray flux . Uniquely, the high yield shots of N210808, N221204, and EYC all demonstrate signs of intense peaks at or after bangtime. It is posited that this is hohlraum wall reheating,rubery where blast-wave-driven detritus from the ablator shell in turn heats the inner surface of the gold-plated, depleted uranium (Au+DU) hohlraum walls. In Table 2 we present a comparison of the metrics for the experimental shots (N180128, N210808, and N221204) relative to our baseline simulations at T. As discussed in Appendix A, the primary goal here was to match closely bangtime and hotspot shape and then to get yield and as close as possible, where is the -th Legendre polynomial and is the burn-weighted ion temperature over space and time. Per standard practices in the ICF community these tunes were quite close, with an exception for N221204 which is about 50% too high in yield. This was deemed acceptable given the rapid rate of performance increase in recent experiments based on the N221204 design. Other metrics of interest are included, such as peak implosion velocity , primary neutron down-scattered ratio (DSR), a figure of merit that is directly proportional to the fuel areal density at stagnation,knauer2 peak radiation temperature , and the burn-up fraction of the DT fuel.
@@
| expbaker | sim | expabu1 | sim | expabu2 | sim | |
@@ III Simulation Methodology and Tuning
All simulations performed in this study were done with the Lasnex rad-MHD code zimmerman75 and in particular via the Lasnex Hohlraum Template (LHT).strozzi1 The LHT is part of an effort to unify standards and physics packages across ICF modeling with Lasnex. All Lasnex simulations were done in RZ-2D geometry and included both the hohlraum and capsule, known as a fully integrated simulation. Lasnex uses 3D ray tracing to propagate lasers through the hohlraum with refraction and inverse bremsstrahlung absorption with corrections due to the Langdon effect.langdon1; sherlock1 NIF has four cones of beams at 23∘, 30∘, 44∘, and 50∘ from the vertical axis. Cross-beam energy transfer (CBET) via ion acoustic waves can transfer energy to the laser with the lower frequency in the plasma flow frame and this is modeled with an inline CBET package in Lasnex.kruer1; michel1; higginson1; strozzi2017
@@
We model non-local thermodynamic equilibrium (NLTE) atomic physics with the Detailed Configuration Accounting (DCA) suite,scott; scott2022 for material properties including opacity and equation of state in hohlraum regions with electron temperatures above 300 eV. We include external LEH hardware, such as an aluminum washer,higginson1 which can affect LEH conditions and CBET. Lasnex also has a robust multi-ion-species hydrodynamics package based on a 13-moment formulation in Ref. schunk and extended MHD that is necessary for this study and generally elevates the fidelity of modeling ICF physics. A more detailed explanation of the MHD package can be found in Ref. strozzi1. In addition, there is a buoyancy-drag, annular shell mix model zhou that we use to degrade simulation yields approximately in line with experimental observations.bachmann This is intended to include the effects of hydrodynamic mix, e.g., due to the Rayleigh-Taylor instability and capsule fill-tube, in simulations that do not resolve this physics. We currently do not have an effect of magnetization on such mix inside our models.
ICF simulations typically overpredict the hohlraum x-ray drive and thereby under-predict capsule bangtime. A common solution to this issue is to use a power multiplier on the drive laser, typically around 0.8 at peak power and additional multipliers in the foot to match shock timings. Opacity multipliers are seen as a more favorable solution, given the higher uncertainty in the NLTE physics versus laser deliver measurements.divol Lower power results in different laser absorption and lower which modifies CBET, but in this study we only used power multipliers since that is a more mature method and the object of our study is not hohlraum physics but rather the implosion itself. We also use tuning parameters for CBET and atomic mix to match implosion shape and fusion yield.
To achieve the best results for matching the simulations to the experiments in question, we go through a multi-step tuning process. In this study, we used the Automated NIF Tuning Suite (ANTS)weber framework for tuning rad-hydro simulations to experiments that utilizes Gaussian-process-based Bayesian optimization. Here, we went through three steps:
@@
Peak power , CBET clamp (the ion-wave electron density fluctuation ), and laser cone fraction multipliers to match bangtime and shape of the 13-15 MeV neutron profile of the primary DT experiment.
@@
A detailed description of the tuning process can be found in Appendix A as well as Ref. schmit.

### v2

@@ Abstract
Motivated by breakthroughs in inertial confinement fusion (ICF), first achieving ignition conditions in National Ignition Facility (NIF) shot N210808 and then laser energy breakeven in N221204,Abu-Shawareb et al. (2022, 2024) modeling efforts here investigate the effect of imposed magnetic fields on integrated hohlraum simulations of igniting systems. Previous NIF experiments have shown fusion yield and hotspot temperature to increase in magnetized, gas-filled capsules. Moody et al. (2022) In this work, we use the 2D radiation-magnetohydrodynamics code LasnexZimmerman and Kruer (1975) with a Livermore ICF common model.Strozzi et al. (2024) Simulations are tuned to closely approximate data from unmagnetized experiments. Investigated here is the effect of imposed axial fields of up to 100 T on the fusion output of high-performing ICF shots, specifically the record BigFoot shot N180128, and HYBRID-E shots N210808 and N221204. The main observed effect is an increase in the hotspot temperature due to magnetic insulation. Namely, electron heat flow is constrained perpendicular to the magnetic field and alpha trajectories transition to gyro-orbits, enhancing energy deposition.Perkins et al. (2017) In addition, we investigate the impact of applied magnetic fields to future NIF designs, specifically an example Enhanced Yield Capability design with 3 MJ of laser energy and a high-, low implosion velocity Pushered-Single-Shell design. In conclusion, magnetization with field strengths of 5-75 T is found to increase the burn-averaged ion temperature by 50% and the neutron yield by 2-12. Specifically, we see yield enhancement of at least 50% with a 5-10 T applied magnetic field for N221204, while a 65 T field on N210808 with drive symmetrization gives an 8 increase in yield.
@@ I Introduction
Magnetized ICF is a relatively old concept, dating back at least to the 1980’s in established literature.Lindemuth and Kirkpatrick (1983); Jones and Mead (1986); Gus’kov et al. (1984) The basic promise is that of boosted ignition conditions and yields via reduced conduction losses and enhanced alpha energy deposition into the ignited hotspot due to enhanced magnetic field strengths, where Tesla-scale applied magnetic fields are compressed by the implosion. In Ref. Lindemuth and Kirkpatrick, 1983, reduced order models informed a relatively large parameter scan suggesting magnetization, starting from an already compressed capsule, could enable designs with less stringent fuel densities, lower implosion velocities, and lower drive requirements that also have significant fuel burn-up, a motivation that continues in recent studies.Ho et al. (2024) Ref. Jones and Mead, 1986 suggests that magnetization can enhance the yield for volumetric burn, particularly when burn propagation does not inherently occur but may be detrimental in robustly burning circumstances. This is the origin of the concept that magnetic fields may inhibit burn propagation in igniting targets. One of the primary conclusions of the current study is that applied magnetic fields do not inherently degrade igniting targets and, in fact, may more than double the yield of such designs.
@@
Experiments have subsequently validated several of the basic ideas of these early modeling studies, with the notable exception that no magnetized, cryogenic implosions with ice layers have been conducted thus far. Cylindrical and spherical implosions with direct laser drive at the OMEGA laser facility Knauer et al. (2010); Chang et al. (2011); Hohenberger et al. (2012) demonstrated field compression and magnetized hotspot insulation according to radiation-magnetohydrodynamic (rad-MHD) modeling. Pulsed-power-driven implosions of magnetized cylindrical liners at the Sandia Z Machine provide similar validation in a different regime, relevant to the MagLIF fusion concept.Slutz and Vesey (2012); Gomez et al. (2014) More recent hohlraum-driven NIF experiments Moody et al. (2022); Sio et al. (2023); Strozzi et al. (2024) with an imposed initial field up to 28 T have shown yield increases up to 2.9x and burn-averaged increases up to 1.4x, consistent with 2D Lasnex rad-MHD modeling.Strozzi et al. (2024) These experiments used gas-filled capsules and 1 MJ of laser energy, due to current practical constraints that magnetized NIF shots must be done at room temperature and with low enough laser energy to avoid risk of optics damage from Brillouin backscatter. Magnetized NIF experiments also examine other topics, like laser-heated gas pipes relevant to MagLIFPollock et al. (2023) and high temperature X-ray sources.Kemp et al. (2016); Leal et al. (2025)
@@
This study is the first to present state-of-the-art, fully-integrated rad-MHD modeling of hohlraum-driven, ignition-class implosions with an imposed magnetic field. Where early modeling studies were restricted to reduced-order rad-MHD calculations with strictly azimuthal fields or capsule-only simulations with a frequency-dependent-source (FDS) drive, we are now able to rigorously model magnetized ICF at an appreciable level of fidelity. Modern rad-MHD codes can model full 2D and 3D capsule and hohlraum systems with expanded, non-ideal MHD. These have found that, with the application of more realistic axial fields, initially 10 T fields can be compressed to 10 kT levels to enhance ICF designs, due to factors such as increased robustness to surface roughness, low-mode shape, and residual kinetic energy.Perkins et al. (2013, 2017); Strozzi et al. (2015); Zimmerman et al. (2014) 3D modeling with the Gorgon code of the planar magnetized Rayleigh-Taylor instability has shown reduced growth for modes with k-vectors parallel to the field,Walsh et al. (2022) though magnetized instability modeling of a spherical, integrated implosion has not been published thus far. In addition, there is work looking into the effect of magnetization on hot electron dynamicsStrozzi et al. (2015) and burn propagation, where some inhibition of burn across field lines has been observed in ignition-style designs.O’Neill et al. (2025)
@@
In this study we model various hohlraum-driven, ignition-class NIF designs and the effect magnetization has on them. It is found that we are able to routinely enhance the yield of full-scale shots like N180128, N210808, and N221204 by at least a factor of while potentially approaching for N210808 at 65 T. The shot nomenclature stands for NIF shot , year , month , and day . There is a notable increase in all three cases with the application of just a 5-10 T magnetic field. The currently studied Enhanced Yield Capability (EYC) design very robustly ignites,MacLaren et al. (2024) meaning the benefit of magnetization is significantly reduced at least for the ad-hoc application of an external field. The Pushered Single Shell (PSS) suggests a way to better take advantage of magnetization given its cooler baseline temperature.MacLaren et al. (2021, 2025); Dewald et al. (2022); Ho et al. (2018)
@@ II NIF Designs for Magnetization
In this study, we focused on five NIF designs for magnetization, three historically high-performing shots, and two potential future designs. The design parameters can be found in Table 1. The selection of shots was not arbitrary; they were all high-performing designs either of historical importance or notable promise based on best practices. All targets used a deuterium-tritium (DT) ice fuel layer and a high-density carbon ablator (HDC) doped with tungsten (W), except for PSS, which used a beryllium (Be) ablator doped with molybdenum (Mo) and traces of argon (Ar) from production. The PSS design is even more unique in that at the outer radius it begins with just Be but then smoothly grades up in density using a dopant. is the difference in wavelength between the inner (23∘, 30∘) and outer (44∘, 50∘) laser cones, which is known as two-color tuning. However, for EYC and PSS we used three- and four-color tuning as well, i.e., separate for 23-30∘ and 44-50∘ quad pairs, to help with shape and reduce potential stimulated Brilloiun scattering (SBS). In this paper we will define in terms of a triplet, where the first number defines two-color wavelength difference (inners versus outers), the second the three-color (23-30∘), and the third the four-color (44-50∘), i.e., but . It is acknowledged that there are many other designs that could be explored, the potential parameter space is enormous. Alternative approaches that could be integrated into a magnetized ICF design were not considered, such as plastic ablators, fast and shock ignition,Strozzi et al. (2012) and wetted foams, for example.
@@
The first shot in consideration was N180128. This record BigFoot shot achieved a yield of kJ and keV Baker et al. (2020) and was the record yield at the time. Besides its historical importance, one intention for investigating this design was to see if magnetization could take a high-performing, non-igniting design and push it over the so-called ignition cliff. This was not found to be the case for N180128 with no other changes but may be possible with further optimization.Walsh et al. (2025a) The second was N210808,Abu-Shawareb et al. (2022) a HYBRID-E (High Yield Big Radius Implosion Design, the E-th iteration) design, which was the first to achieve greater than 1 MJ yield and exceed all extant Lawson criteria for ICF ( MJ, keV). The third case example was N221204,Abu-Shawareb et al. (2024) the first shot to achieve energy breakeven, where more energy was generated than delivered to the hohlraum ( MJ, keV). There have been many minor, cumulative modifications in designs with each shot, but the most pronounced change had been in the increase in laser energy ( MJ, MJ, MJ), and notably the greater radiation temperature efficiency of the Hybrid-E hohlraum.
@@
Two future designs were considered as well. Specifically, a preliminary EYC design,MacLaren et al. (2024) that extends standard HYBRID-E designs from 2 MJ to 3 MJ of laser energy with a peak power of 450 TW. Simple hydroscaling of a design like N221204 is not an option due to laser optics damage exacerbated by higher laser powers. Therefore, new design parameter spaces must be explored. The second is the PSS design driven by 1.9 MJ of laser energy,MacLaren et al. (2021); Dewald et al. (2022); Ho et al. (2018) which is a platform tailored to radiochemistry studies and uses a density gradient of Mo-doped Be as the ablator instead of CH or HDC. This design is perhaps ideal for magnetization because it inherently has a slower, high , low-temperature implosion that can be directly enhanced by the application of an external magnetic field.
@@
In Fig. 1 we show a depiction of the hohlraum and target capsules relative to the experimental properties in Table 1. Four of the shots had qualitatively similar HDC capsule designs, while PSS has a unique design. A comparison of the laser power profiles and corresponding hohlraum radiation temperature time history can be seen in Figs. 2 (a) and (b), where is the average radiation temperature across a spherical annulus outside the ablation front of the capsule. N180128 has a distinct BigFoot design while N210808, N221204, and EYC all have HYBRID-E designs. PSS has a distinctly longer laser foot because Be does not need a picket, which is deleterious to symmetry control, due its much lower melt pressure that keeps the 1st shock rarefaction pressure above re-freeze and better ablation front stability.Zylstra et al. (2018) HDC on the other hand uses a picket due to the higher melt temperature. As is apparent, HYBRID-E designs presented a more efficient approach than BigFoot while EYC has an extended peak to take advantage of greater energy without the higher peak power dangerous to the optics. In Fig. 2 (b) we have raised to the fourth power and normalized to a standard value of (300 eV)4 to represent x-ray flux . Uniquely, the high yield shots of N210808, N221204, and EYC all demonstrate signs of intense peaks at or after bangtime. It is posited that this is hohlraum wall reheating,Rubery et al. (2024) where blast-wave-driven detritus from the ablator shell in turn heats the inner surface of the gold-plated, depleted uranium (Au+DU) hohlraum walls. In Table 2 we present a comparison of the metrics for the experimental shots (N180128, N210808, and N221204) relative to our baseline simulations at T. As discussed in Appendix A, the primary goal here was to match closely bangtime and hotspot shape and then to get yield and as close as possible, where is the -th Legendre polynomial and is the burn-weighted ion temperature over space and time. Per standard practices in the ICF community these tunes were quite close, with an exception for N221204 which is about 50% too high in yield. This was deemed acceptable given the rapid rate of performance increase in recent experiments based on the N221204 design. Other metrics of interest are included, such as peak implosion velocity , primary neutron down-scattered ratio (DSR), a figure of merit that is directly proportional to the fuel areal density at stagnation,Knauer et al. (2011) peak radiation temperature , and the burn-up fraction of the DT fuel.
@@
| expBaker et al. (2020) | sim | expAbu-Shawareb et al. (2022) | sim | expAbu-Shawareb et al. (2024) | sim | |
@@ III Simulation Methodology and Tuning
All simulations performed in this study were done with the Lasnex rad-MHD code Zimmerman and Kruer (1975) and in particular via the Lasnex Hohlraum Template (LHT).Strozzi et al. (2024) The LHT is part of an effort to unify standards and physics packages across ICF modeling with Lasnex. All Lasnex simulations were done in RZ-2D geometry and included both the hohlraum and capsule, known as a fully integrated simulation. Lasnex uses 3D ray tracing to propagate lasers through the hohlraum with refraction and inverse bremsstrahlung absorption with corrections due to the Langdon effect.Langdon (1980); Sherlock et al. (2024) NIF has four cones of beams at 23∘, 30∘, 44∘, and 50∘ from the vertical axis. Cross-beam energy transfer (CBET) via ion acoustic waves can transfer energy to the laser with the lower frequency in the plasma flow frame and this is modeled with an inline CBET package in Lasnex.Kruer et al. (1996); Michel et al. (2009); Higginson et al. (2022); Strozzi et al. (2017)
@@
We model non-local thermodynamic equilibrium (NLTE) atomic physics with the Detailed Configuration Accounting (DCA) suite,Scott and Hansen (2010); Scott et al. (2022) for material properties including opacity and equation of state in hohlraum regions with electron temperatures above 300 eV. We include external LEH hardware, such as an aluminum washer,Higginson et al. (2022) which can affect LEH conditions and CBET. Lasnex also has a robust multi-ion-species hydrodynamics package based on a 13-moment formulation in Ref. Schunk, 1977 and extended MHD that is necessary for this study and generally elevates the fidelity of modeling ICF physics. A more detailed explanation of the MHD package can be found in Ref. Strozzi et al., 2024. In addition, there is a buoyancy-drag, annular shell mix model Zhou et al. (2002) that we use to degrade simulation yields approximately in line with experimental observations.Bachmann et al. (2023) This is intended to include the effects of hydrodynamic mix, e.g., due to the Rayleigh-Taylor instability and capsule fill-tube, in simulations that do not resolve this physics. We currently do not have an effect of magnetization on such mix inside our models.
ICF simulations typically overpredict the hohlraum x-ray drive and thereby under-predict capsule bangtime. A common solution to this issue is to use a power multiplier on the drive laser, typically around 0.8 at peak power and additional multipliers in the foot to match shock timings. Wall opacity multipliers are seen as a more favorable solution, given the higher uncertainty in the NLTE physics versus laser deliver measurements.Divol et al. (2024) Lower power results in different laser absorption and lower which modifies CBET, but in this study we only used power multipliers since that is a more mature method and the object of our study is not hohlraum physics but rather the implosion itself. We also use tuning parameters for CBET and atomic mix to match implosion shape and fusion yield.
To achieve the best results for matching the simulations to the experiments in question, we go through a multi-step tuning process. In this study, we used the Automated NIF Tuning Suite (ANTS)Weber et al. (2024) framework for tuning rad-hydro simulations to experiments that utilizes Gaussian-process-based Bayesian optimization. Here, we went through three steps:
@@
Peak power , CBET clamp (the ion-wave electron density fluctuation ), and laser cone fraction multipliers to match bangtime and shape of the 13-15 MeV neutron profile of the primary DT experiment. The CBET clamp defaults to 0.005 before being changed later in time after the laser foot, while the CF multiplier when used is constant throughout.
@@
A detailed description of the tuning process can be found in Appendix A as well as Ref. Schmit et al., 2025.